OSCILLATOR STRENGTHS AND PREDISSOCIATION RATES FOR RYDBERG COMPLEXES IN ^{12}C^{16}O BETWEEN 92.9 AND 93.4 NM

M. EIDELSBERG, J. L. LEMAIRE, Observatoire de Paris, Paris, France; \underline{S. R. FEDERMAN}, Department of Physics and Astronomy, University of Toledo, Toledo, OH 43606; G. STARK, A. N. HEAYS, Department of Physics, Wellesley College, Wellesley, MA 02481; L. GAVILAN, Observatoire de Paris, Paris, France; J. H. FILLION, Université PVI UMPC, Paris, France; F. ROS-TAS, Observatoire de Paris, Paris, France; J. R. LYONS, IGPP, University of Califormia, Los Angeles, CA 90095; P. L. SMITH, Department of Physics, Wellesley College, Wellesley, MA 02481; N. DE OLIVEIRA, D. JOYEUX, M. ROUDJANE, L. NAHON, Synchrotron SOLEIL, Saint Aubin, France.

We are conducting experiments on the DESIRS beam-line at the SOLEIL Synchrotron to acquire the necessary data on oscillator strengths and predissociation rates for modeling CO photochemistry in astronomical environments. A VUV Fourier Transform Spectrometer provides a resolving power of about 350,000, allowing us to discern individual lines in electronic transitions. Here we focus on results for six overlapping bands seen in spectra of ^{12}C^{16}O and compare them with earlier determinations. Absorption from the ground electronic level, X ^{1}Σ^{+} v′′ = 0, to the upper levels 4pπ(2), II 1Π, 4pσ(2), 5pπ(0), 5pσ(0), and I ^{1}Π was analyzed. Our results are the first to provide data on each band. The spectra also reveal a continuum feature that was modeled, but its identification is not known at the present time.